# AccDiffusion v2: Towards More Accurate Higher-Resolution Diffusion Extrapolation

### v1

@@ 1 Introduction
Our contributions are summarized as follows: (1) We identify the reason for repetitive generation during patch-wise denoising and introduce patch-content-aware prompts to effectively suppress this issue. (2) We uncover the cause of local distortion during patch-wise denoising and incorporate low-resolution structure information into the patch-wise denoising process using ControlNet to effectively suppress it. (3) We propose dilated sampling with interaction to generate more accurate global semantic information, effectively reducing both repetitive generation and local distortion. (4) We conduct a thorough comparison with the latest higher-resolution image generation methods, demonstrating that our approach achieves state-of-the-art performance in training-free image generation extrapolation.
A preliminary version of this paper, termed AccDiffusion, can be referred to the publication [24]. Building on that version, this paper further explores the cause of local distortion in higher-resolution image generation and introduces AccDiffusion v2 to suppress it effectively. We also discuss the pros and cons of the latest related works in this field and provide a more comprehensive comparisons to highlight the advantages of AccDiffusion v2. Beyond these changes, we conduct a more comprehensive ablation study and failure case analysis in this paper. This study, while providing valuable insights, is not without its shortcomings. Therefore, we also identify the limitations of this paper and suggest potential directions for future research.
@@ 4.1 Patch-Content-Aware Prompts
While DemoFusion showcases the potential of leveraging pre-trained LDMs to generate higher-resolution images, the persistent issue of small object repetition poses a challenge to its performance, as depicted in Fig. 1(d). To pinpoint the cause of this repetition, we design two ablation experiments. In the first one, we exclude the text prompt during higher-resolution generation of DemoFusion. The result in Fig. 4(a) shows that the removal of prompts completely eliminates repetitive objects but results in a noticeable loss of detail; In the second one, we exclude the operations of residual connection & dilated sampling in DemoFusion. The result in Fig. 4(b) suffers severe large object repetition. From these results, it is reasonable to conclude that small object repetition arises as an adverse effect from using the same text prompt across all patches, as well as from residual connection and dilated sampling operations. While the former promotes object repetition, the latter diminishes it. As a result, DemoFusion tends to generate small repetitive objects.
The above analysis reveals that simply excluding text prompts during higher-resolution generation to eliminate small object repetition is not a feasible remedy, as it would inevitably result in a compromise on image fidelity. Considering the significant role that text prompts play in image generation and the inaccuracy of identical text prompts for all patches, it is crucial to tailor more accurate prompts for each patch. That is, if an object is not present in a patch, the corresponding word in the text prompts should not serve as a prompt for that patch.
Figure 4: Results of higher-resolution image generation. (a) The result of DemoFusion without text prompt. (b)The result of DemoFusion without residual connection and dilated sampling. (c) The result of dilated sampling without window interaction. (d)The result of our dilated sampling with window interaction. Best viewed by zooming in.
@@
where $j$ and $i$ enumerates $M$ and $P_{1}$, respectively. The hyper-parameter $c\in(0,1)$ determines if the proportion of a highly responsive region corresponding to a word $y_{j}$ exceeds the threshold for inclusion in the prompts of patch $z_{t}^{i}$. We then concatenate all words that should appear in a patch together, resulting in patch-content-aware prompts $\{\gamma^{i}\}_{i=1}^{P_{1}}$ for noise patches $\{z_{t}^{i}\}_{i=1}^{P_{1}}$ during patch-wise denoising.
@@ 4.3 Dilated Sampling with Window Interaction
where $(G_{t-1})_{i::h_{s},j::w_{s},:}={\mathcal{D}^{k}_{t-1}}$ and $\eta$ decreases from $1$ to 00 following a cosine schedule. As shown in Fig. 4(c), we find that the global semantic information is non-smooth, due to the lack of interaction among different samples. To solve this issue, as illustrated in Fig. 7, we enable window interaction among different samples prior to each denoising process through a bijective function:
@@
where $f^{h,w}_{t}$ is a bijective function, with the mapping varying on the specific position or time step. We then perform standard denoising progress on $\{\mathcal{D}_{t}^{k}\}_{k=1}^{P_{2}}$ to obtain $\{\mathcal{D}_{t-1}^{k}\}_{k=1}^{P_{2}}$. Before applying Eq. (15) to $\{\mathcal{D}_{t-1}^{k}\}_{k=1}^{P_{2}}$, we recover the position by using the inverse mapping ${(f^{h,w}_{t})}^{-1}$ of $f^{h,w}_{t}$ as:
@@
which yields more smooth global semantics like Fig. 4(d).
@@ 5.2 Quantitative Comparison
We employ three widely-used metrics: Frechet Inception Distance (FID) [40], Inception Score (IS) [41], and CLIP Score [42] for quantitative evaluations. Specifically, $\text{FID}_{r}$ assesses the Frechet Inception Distance between generated high-resolution images and real images, while $\text{IS}_{r}$ calculates the Inception Score for these generated high-resolution images. Notably, both $\text{FID}_{r}$ and $\text{IS}_{r}$ require resizing images to $299^{2}$ resolutions, which may not provide optimal assessments for high-resolution images. To address this, inspired by methods  [11, 43], we crop 10 local patches at native resolution (1x) from each generated high-resolution image before resizing, yielding $\text{FID}_{c}$ and $\text{IS}_{c}$. The CLIP Score is calculated based on the cosine similarity between image embeddings and text prompts, providing an additional alignment metric. For quantitative comparison, we randomly selected $10,000$ images from the Laion-5B dataset [44] as the real image set and used $1,000$ randomly chosen text prompts from Laion-5B as input for AccDiffusion v2, generating a corresponding set of high-resolution images.
@@ 5.5 Ablation Study
5.5.1 Ablations on Core Modules.
Fig. 10 illustrates that removing any module reduces generation quality. Excluding patch-content-aware prompts leads to numerous small, repetitive object repetitions, emphasizing the role of patch-content-aware prompts in preventing repetitive generation. When dilated sampling with window interaction is removed, small objects in the image appear unrelated to the image, demonstrating that dilated sampling with window interaction enhances semantic consistency and minimizes repetition. Moreover, without ControlNet-assisted generation, the image exhibits local distortion, indicating that ControlNet helps establish more accurate local structures. When all modules are removed, the image shows the most repetitive objects; however, using all modules together effectively prevents both repetitions and local distortion. This demonstrates that these modules function collectively to minimize artifacts.
@@
5.5.2 Ablations on Hyper-Parameters.
@@ 6 Limitations and Future work
While providing valuable insights, AccDiffusion v2 is not without its shortcomings: Firstly, the inference latency of AccDiffusion v2 is high as shown in Table I, akin to other patch-wise extrapolation methods [24, 11], due to inefficient progressive upscaling and overlapped patch-wise denoising. Additionally, the use of ControlNet to suppress local distortion further adds to this delay. However, users can choose the method that best fits their needs with this trade-off between performance and inference latency. Secondly, the fidelity of extrapolation results heavily relies on the pre-trained diffusion model, given that AccDiffusion v2 is training-free. Consequently, stronger diffusion models lead to improved AccDiffusion v2 performance and vice versa. Thirdly, as depicted in Fig. 13, AccDiffusion v2 excels in generating intricate details like the cat’s eye but may introduce irrelevant elements such as superfluous “PPP” details. Lastly, both AccDiffusion and patch-wise methods should allow infinite extrapolation. However, when the resolution exceeds 8K (64×), AccDiffusion v2, along with existing techniques [11, 24], encounters detail degradation.
To enhance efficiency, forthcoming research could explore non-overlapping patch-based denoising techniques to alleviate inference latency. AccDiffusion v2 offers valuable insights, decoupling the image-content-aware prompt into patch-content-aware prompts to address object repetition caused by inaccurate prompts. Future works have an opportunity to use vision large language models [45] or image caption models [46] to refine prompts for distinct patches. Furthermore, AccDiffusion v2 highlights the benefits of incorporating extra controls like ControlNet [23] to generate coherent high-resolution image structures. Future studies could delve into employing controllable generation methods for high-resolution image extrapolation.
@@ Acknowledgments
This work was supported by National Science and Technology Major Project (No. 2022ZD0118202), the National Science Fund for Distinguished Young Scholars (No.62025603), the National Natural Science Foundation of China (No. U21B2037, No. U22B2051, No. U23A20383, No. 62176222, No. 62176223, No. 62176226, No. 62072386, No. 62072387, No. 62072389, No. 62002305 and No. 62272401), and the Natural Science Foundation of Fujian Province of China (No.2022J06001).

### v2

@@ 1 Introduction
The conference version of this paper, termed AccDiffusion, is published in [24]. In the conference version, our contributions are summarized as follows: (1) We identify the reason for repetitive generation during patch-wise denoising and introduce patch-content-aware prompts to effectively suppress this issue. (2) We propose dilated sampling with interaction to generate more accurate global semantic information, effectively reducing repetitive generation.
In this journal version, we build upon the conference version and introduce several substantial improvements as follows: (1) We conduct a more in-depth analysis of the causes of local distortion during patch-wise denoising. To address this issue, we incorporate low-resolution structural information into the patch-wise denoising process using ControlNet, effectively mitigating local distortions and addressing the limitations raised in the conference version. (2) We perform more extensive experiments, both quantitatively and qualitatively. Quantitatively, we compare AccDiffusion v2 with the latest related methods, achieving updated state-of-the-art performance. Qualitatively, we provide comparisons with more related methods (Fig. 9) and showcase additional results at various resolutions (Fig. 10). Furthermore, we reconduct ablation studies (Sec. 5.5.1) to validate the compatibility of our new module with those proposed in the conference version. We also apply AccDiffusion v2 to different stable diffusion variants (Fig. 11). (3) While this study provides valuable insights, it is not without its shortcomings. Therefore, we include an analysis of failure cases (Fig. 15), identify the limitations of this work, and suggest potential directions for future research.
@@ 4.1 Patch-Content-Aware Prompts
Current image generation extrapolation methods [10, 11] suffer from repetitive generation as shown in Fig. 1 and Fig. 4. This repetition typically falls into two categories: large object repetition and small object repetition. For example, MultiDiffusion [10], which use the same prompts across all patches, generates multiple large repeated objects scattered across the image. By integrating residual connection & dilated sampling, DemoFusion [11] addresses large object repetition but leads to small object repetition in the background as shown in Fig. 4(b). To identify the cause of small object repetition, we perform an ablation experiment by excluding the text prompt during higher-resolution generation of DemoFusion. The result in Fig. 4(c) shows that the removal of prompts completely eliminates small repetitive objects but results in a noticeable loss of detail. From these results, it is reasonable to conclude that small object repetition arises as an adverse effect from using the same text prompt across all patches, as well as from residual connection and dilated sampling operations. While the former promotes object repetition (Fig. 4(a)), the latter diminishes it (Fig. 4(b)). As a result, DemoFusion tends to generate small repetitive objects. The above analysis reveals that simply excluding text prompts during higher-resolution generation to eliminate small object repetition is not a feasible remedy, as it would inevitably result in a compromise on image fidelity. Considering the significant role that text prompts play in image generation and the inaccuracy of identical text prompts for all patches, it is crucial to tailor more accurate prompts for each patch. That is, if an object is not present in a patch, the corresponding word in the text prompts should not serve as a prompt for that patch.
Figure 4: Results of higher-resolution image generation. (a) MultiDiffusion suffers from large object repetition. (b) DemoFusion suffers from small object repetition. (c) DemoFusion without text prompts results in a severe loss of details. (d) Our method, utilizing patch-content-aware prompts, effectively suppresses small object repetition while preserving image fidelity. Best viewed by zooming in.
@@
where $j$ and $i$ enumerates $M$ and $P_{1}$, respectively. The hyper-parameter $c\in(0,1)$ determines if the proportion of a highly responsive region corresponding to a word $y_{j}$ exceeds the threshold for inclusion in the prompts of patch $z_{t}^{i}$. We then concatenate all words that should appear in a patch together, resulting in patch-content-aware prompts $\{\gamma^{i}\}_{i=1}^{P_{1}}$ for noise patches $\{z_{t}^{i}\}_{i=1}^{P_{1}}$ during patch-wise denoising. Patch-content-aware prompts effectively suppress small object repetition while preserving image fidelity (Fig. 4(d)).
@@ 4.2 More Accurate Generation of Local Content
Figure 7: Results of 4×4$\times$4 × higher-resolution image generation. (a) The result of dilated sampling without window interaction. (b)The result of our dilated sampling with window interaction. Best viewed by zooming in.
@@ 4.3 Dilated Sampling with Window Interaction
where $(G_{t-1})_{i::h_{s},j::w_{s},:}={\mathcal{D}^{k}_{t-1}}$ and $\eta$ decreases from $1$ to 00 following a cosine schedule. As shown in Fig. 7(a), the global semantic information generated by dilated sampling appears inconsistent and non-smooth. This issue arises because dilated sampling independently denoises $\{D_{t}^{k}\}_{k=1}^{P_{2}}$ and subsequently combines the denoised outputs $\{\mathcal{D}_{t-1}^{k}\}_{k=1}^{P_{2}}$ to form the global semantic information. The lack of interaction among different samples $\mathcal{D}_{t}^{k}$, as illustrated in Fig. 7(a), results in discrepancies between the denoised outputs, leading to non-smooth global semantic information. To solve this issue, as illustrated in Fig. 8, we enable window interaction among different samples prior to each denoising process through a bijective function:
@@
where $f^{h,w}_{t}$ is a bijective function with random mapping, with the mapping varying on the specific position or time step. We then perform standard denoising progress on $\{\mathcal{D}_{t}^{k}\}_{k=1}^{P_{2}}$ to obtain $\{\mathcal{D}_{t-1}^{k}\}_{k=1}^{P_{2}}$. Before applying Eq. (15) to $\{\mathcal{D}_{t-1}^{k}\}_{k=1}^{P_{2}}$, we recover the position by using the inverse mapping ${(f^{h,w}_{t})}^{-1}$ of $f^{h,w}_{t}$ as:
@@
which yields more smooth global semantics like Fig. 7(b). This verifies that through window interaction, different samples $\{D_{t}^{k}\}_{k=1}^{P_{2}}$ can exchange information with each other, which is crucial for generating more consistent and smoother global semantic information.
@@ 5.2 Quantitative Comparison
We employ three widely-used metrics: Frechet Inception Distance (FID) [40], Inception Score (IS) [41], and CLIP Score [42] for quantitative evaluations. Specifically, $\text{FID}_{r}$ assesses the Frechet Inception Distance between generated high-resolution images and real images, while $\text{IS}_{r}$ calculates the Inception Score for these generated high-resolution images. Notably, both $\text{FID}_{r}$ and $\text{IS}_{r}$ require resizing images to $299^{2}$ resolutions, which may not provide optimal assessments for high-resolution images. To address this, inspired by methods  [11, 43], we crop 10 local patches at native resolution (1×1$\times$1 ×) from each generated high-resolution image before resizing, yielding $\text{FID}_{c}$ and $\text{IS}_{c}$. The CLIP Score is calculated based on the cosine similarity between image embeddings and text prompts, providing an additional alignment metric. For quantitative comparison, we randomly selected $10,000$ images from the Laion-5B dataset [44] as the real image set and used $1,000$ randomly chosen text prompts from Laion-5B as input for AccDiffusion v2, generating a corresponding set of high-resolution images.
@@ 5.5 Ablation Study
5.5.1 Ablations on Core Modules
Fig. 12 illustrates that removing any module reduces generation quality. Excluding patch-content-aware prompts leads to numerous small, repetitive object repetitions, emphasizing the role of patch-content-aware prompts in preventing repetitive generation. When dilated sampling with window interaction is removed, small objects in the image appear unrelated to the image, demonstrating that dilated sampling with window interaction enhances semantic consistency and minimizes repetition. Removing dilated sampling with window interaction leads to small object occurrences in parts of the big object, e.g., a small cat head on the claw of the big cat. This indicates that dilated sampling with window interaction helps maintain the overall structure of the image by providing accurate global structure information. Moreover, without ControlNet-assisted generation, the image exhibits local distortion, indicating that ControlNet helps establish more accurate local structures. When all modules are removed, the image shows the most repetitive objects; however, using all modules together effectively prevents both repetitions and local distortion. This demonstrates that these modules function collectively to minimize artifacts.
@@ 5.5.1 Ablations on Core Modules
5.5.2 Ablations on Hyper-Parameters
@@
Figure 14: Ablation study on the number of ControlNet steps. Local distortions are highlighted with red boxes. The inference time is measured on one NVIDIA RTX 3090 GPU. The prompt used for the image is “A cat is playing with furry toys on the lawn.” Best viewed by zooming in.
5.5.3 Ablations on ControlNet Integration
In Sec. 4.2, we integrate ControlNet into our framework to enhance the fidelity of local structures during higher-resolution extrapolation. While this integration introduces additional computational overhead, it suppresses local artifacts in the results. To rigorously evaluate the trade-off between image quality and inference efficiency, we perform an ablation study by varying the number of ControlNet steps. Specifically, we alternate between AccDiffusion v2 without ControlNet and AccDiffusion v2 with ControlNet during denoising, allowing us to flexibly adjust the frequency of ControlNet invocation. For example, when the ControlNet steps are set to 25, it is applied once every two steps. As shown in Fig. 14, increasing the number of ControlNet steps yields notable improvements in image quality, primarily through the suppression of local artifacts. However, this enhancement comes at the expense of increased inference time. Accordingly, the number of ControlNet steps can be adjusted to balance perceptual quality and inference efficiency based on application requirements.
The additional computational cost in AccDiffusion v2 arises from the inclusion of ControlNet’s condition encoder. To mitigate this, one potential solution is the adoption of lightweight alternatives such as ControlNeXt [45], which maintains comparable guidance performance with reduced latency. Moreover, on high-end GPUs equipped with enhanced parallelism (e.g., NVIDIA H100), the relative overhead introduced by ControlNet is significantly alleviated. For instance, in the 16×16$\times$16 × extrapolation setting, AccDiffusion v2 completes generation in 11 minutes on an H100 GPU, compared to 9 minutes for AccDiffusion.
@@ 6 Limitations and Future work
While providing valuable insights, AccDiffusion v2 is not without its shortcomings: Firstly, the inference latency of AccDiffusion v2 is high as shown in Table I, akin to other patch-wise extrapolation methods [24, 11], due to inefficient progressive upscaling and overlapped patch-wise denoising. Additionally, the use of ControlNet to suppress local distortion further adds to this delay. Secondly, the fidelity of extrapolation results heavily relies on the pre-trained diffusion model, given that AccDiffusion v2 is training-free. Consequently, stronger diffusion models lead to improved AccDiffusion v2 performance and vice versa. Thirdly, as depicted in Fig. 15, AccDiffusion v2 excels in generating intricate details like the cat’s eye but may introduce irrelevant elements such as superfluous “PPP” details. Lastly, both AccDiffusion and patch-wise methods should allow infinite extrapolation. However, when the resolution exceeds 8K (64×), AccDiffusion v2, along with existing techniques [11, 24], encounters detail degradation.
To enhance efficiency, forthcoming research could explore non-overlapping patch-based denoising techniques and use a lightweight condition encoder to alleviate inference latency. AccDiffusion v2 offers valuable insights, decoupling the image-content-aware prompt into patch-content-aware prompts to address object repetition caused by inaccurate prompts. Future works have an opportunity to use vision large language models [46] or image caption models [47] to refine prompts for distinct patches. Furthermore, AccDiffusion v2 highlights the benefits of incorporating extra controls like ControlNet [23] to generate coherent high-resolution image structures. Future studies could delve into employing controllable generation methods for high-resolution image extrapolation.
@@ Acknowledgments
This work was supported by the National Science Fund for Distinguished Young Scholars (No.62025603), the National Natural Science Foundation of China (No. U21B2037, No. U22B2051, No. U23A20383, No. 62176222, No. 62176223, No. 62176226, No. 62072386, No. 62072387, No. 62072389, No. 62002305 and No. 62272401), and the Natural Science Foundation of Fujian Province of China (No. 2021J06003, No.2022J06001).